OpenStudy (anonymous):

How do graphs of y=1/x and y=3/x-4 compare

OpenStudy (mathmale):

That factor of 3 in the 2nd equation will stretch the graph of the first equation vertically by a factor of 3. Try graphing y=1/x, and then y=3/x. That (x+4) in the denominator (my assumption) of the 2nd function will shift the entire graph of y=3/x four units to the left. Experiment with this. Use the Draw function, below. Once I've seen your work, I'd be glad to give you more pointers.